You want to buy a Disney bond with two years until maturity and a coupon rate of 9.50% per year, paid semiannually. FED reports that the market interest rate for similar bonds is only 3.8% per half-year. a. Find the bond's price today and 6 months from now after the next coupon is paid. (Do not round intermediate calculations.

Round your answers to 2 decimal places.)

-Current price $

-Price after 6 months $

- What is the total rate of return on the bond?

(Do not round intermediate calculations. Round your answer to 2 decimal places.) Total rate of return % per six months
Answer #1

n=2*2=4
Coupon =0.095*1000/2=47.50
rate = 0.038
Using excel:
=PV(0.038,4,47.5,1000)
=$1,034.65

Price after 6 months:

n=3
Coupon =0.095*1000/2=47.50
rate = 0.038
=PV(0.038,3,47.5,1000)
=$1,026.46

Capital gain return:=1026.46/1034.65=-0.791%

Coupon return = 47.5/1034.65
=+4.591%

Total return = 3.80%